Obtain the largest possible domain, and corresponding range, of the function f, defind by f(x)=x-2/3-x
1
Expert's answer
2020-05-18T09:26:23-0400

Given function is "f(x)=\frac{x-2}{3-x}" .

Now given function has numerator and denominator both continuous and denominator is "3-x".

So given function is not defined when "3-x=0 \implies x=3".

So, posible domain where function is defined should not contains x=3.

Hence Largest possible domain is D = "(-\infin , 3) \cup (3,\infin)."


Now when "x\in (-\infin,3) \implies x<3 \implies 3-x>0" and when "x\in (3,\infin) \implies 3-x<0."

Also, for "x\in (3,\infin), x-2\in (1,\infin)" and so "f(x)\in (-\infin,-1)."

And for "x\in (-\infin, 3), f(x)=-1+\frac{1}{3-x} \in -1+(0,\infin) \in (-1,\infin)."

So, Range of given function is "(-\infin,-1)\cup(-1,\infin)."